Showing posts with label STA301. Show all posts

STA301 Current Assignment Solution No. 1 Fall 2012 [us]

STA301 Assignment No.1 Last Date: 7th November, 2012
a) Following are the heights (in centimeters) of 45 female high school students. Prepare a frequency distribution of the data where the class width is 5, the number of classes is 6 and the smallest observation is 148. Also obtain the relative frequency and relative cumulative frequency.
 170, 151, 154, 160, 158, 154, 171, 156, 160, 157, 160, 157, 148, 165, 158, 159, 155, 151, 152, 161, 156, 164, 156, 163, 174, 153, 170, 149, 166, 154, 166, 160, 160, 161, 154, 163, 164, 160, 148, 162, 167, 165, 158, 158, 176.
b) Which level of measurement (scale of measurement) is suitable for the following data in each example?
i) The lake Saif-ul-Muluk is 10,578 feet above from the sea level.
ii) If customers go to the shop for buying shoes and try different size of shoes. The sizes of shoes are 7B, 10D and 12E.
iii) The numbering on T-shirts of players in the cricket team.
iv) The weight of the person is 68 kg.

Question 2: Marks: 5+3=8
a) Find the median, harmonic mean and 23rd percentile for the following frequency distribution.
Class Boundaries Frequency
145 – 150 4
150 – 155 6
155 – 160 28
160 – 165 58
165 – 170 64
170 – 175 30

b) Explain the concept of Geometric mean? Also find the geometric mean for the following data.
18, 19, 19, 19, 19, 20, 20, 20, 20, 20, 21, 21, 21, 21, 22, 23, 24, 27, 30, 36
Solution

STA301 VU Current Final Paper Spring 2012

Mcqs=20 marks too much easy and 4rom old papers
subjective questions
1.what s b!as?(2)
2.advantages and disadvantages of medan?(3)
3. Mathematical expecton of discrete random variable?(3)
4.any two properties of mathematical expectation?(2)
5.what s stat!c test?(2)
6. Decide a small sample and large sample?(2)
7.4 questions were numeric 4rom probablty and bnomal dstrbuton?(2 or 3 marks)
8. Calculate the class boundares,class marks and relatve frquencey?(5)
9.2 questions k smagh na aye pata nah kasy kernay thay (5)
10.tell the null and alternative hypothesis of 150?(2)

STA301 Assignment Idea Solution june 2012

Answer No .1

a)      The probability density function of a random variable is given as
F(x) = 2(x-1)               1<x<2
0                                        Elsewhere
      Calculate second moment about origin=?
      (As we know the moment about origin)
       u1 = E (x) ……… first moment
       u2 = E (x) ………. second moment
          u3 = E (x)…………third moment   
       u4 = E (x)…………fourth moment

        As per rave
        P (z≥-1.1)
        -1.1 se bare

      In the second diagram 0.5+0.5=1
      Shaded area is raga

     In the third diagram
     P (z≥ -1.1) =0.5+0.3643
     =0.8643 Answer  

   Second Moment about Origin

   Second moment about origin = E (x²­)

  E (x²­) = -∞∫∞ x2.   f(x) dx
 
  E (x²­) = 1∫2x2. 2 (x+1) dx

  E (x²­) = 21∫2 x2. (x-1) dx

  E (x²­) = 21∫2 (x3 – x2) dx

  E (x²­) = 2 [x4/5 – x3/3]2

  E (x²­) = 2 [(2)4/4 – (2)3/3   -   ( (1)4/4 – (1)3/3 ) ]

  E (x²­) = 2 (48-32/12) - (3-9/12)

   E (x²­) = 2 [16/12 + 1/12]
      

   E (x²­) = 2 (17/12)


 E (x²­) = 17/6 Answer

 Part b)
 
 F (x,y) = x (1+3y2) / 4

 = 0 , elsewhere

    0 < x < 2 , 0 < y < 1
Find the marginal probability density function of x =?

 g(x) = -∞∫∞ f (x,y) dy

 0 < y < 1

   = 0∫1 x (1+ 3y2) /4 .dy

   = 0 ∫1 x (1+ 3y2) /4 .dy

   = 1/4 0∫1 (x + 3xy2) dy
                                   Y=1
   = 1/4 [ xy+ 3xy3 /3]10

    = 1/4 [xy +xy3]1

    =1/4 [x(1) + x(1)3 – (x(0) + x(0)3]

   =1/4 [x+x-0]

   =1/4 (2x)

G(x) = x/2       for (0 < x < 2)

Answer No.2

a)      what is the probability that a poker hand of 5 cards contain exactly 2 Aces ?

n (S) = 52C5

Let a donate exactly 2 aces



             (4C2)      (48C3)
P(A) = ______________
              (52C5)

P(A)=  6 (17296)
            2598960
         

           103776
P(A)= ______         Answer
           2598960 

Part b)
A pair of dice is rolled 180 times. By using the normal approximation to binominal distribution , find the probability of a total of 7 occurs at least 25 times ?

Probability of 7 is 

(1,6) (2,5) (3,4) (4,3) (5,2) (6,1)

P = 6/36 = 1/36

Now

n=180
p= 1/6
q= 1-p = 1-1/6  = 5/6

µ = n p =180*1/6 =30

 ♪ = √ n p q

 ♪ = √ 180*1/6*5/6

 ♪ = 5

Now

Z = X - µ
         ♪

Z= X-30
       5
At least 25 times
Which means

P (x ≥ 25)
Continuity correction

                         Working Notes
when we go to continuous form discrete we add or less 0.5

In this case he said at least 25 times which means 25 is include that is why we less 0.5 from 25 for continuity correction.

P ( x ≥ 24.5)
Now

Z = X - µ
         ♪
Z = X-30
         5

Z = 24.5 – 30
            5
Z = -1.1 

Now
Go to page # 225 of Sta301 handouts and see the value of 1.1 in 0 column.

Because it is 1.10 last pay zero ha agar 1.12 rate ho to 1.1 ki value column 2 me dekhein.
Now 1.1   in column zero = 0.3643
Is ka matlab ha k origin say -1.1 ka fasla 0.3643 ha.

Related Posts Plugin for WordPress, Blogger...